OpenStudy (blenchiki):

Can someone help on how to solve this? Find the value of x. If necessary, round to the nearest tenth

OpenStudy (blenchiki):



OpenStudy (sanoofy):

A=bh/2 A= (9.89)(9.89) or 98 squared/2 A=97.8121/2 A=48.9 Rounded: 49 sq. in (btw, dont forget units) Therefore x=9.8

OpenStudy (blenchiki):

thank you so much! and thank you for explaining how lol @sanoofy

OpenStudy (sanoofy):

No problem! Anytime!